what is the y intercept of x-2y=0
Found 3 solutions by jeanevyazuelo, solver91311, lenny460:
Answer by jeanevyazuelo(1)

let (x=0)
x-2(0)=0
x=0
odered pair (0,0)

Answer by solver91311(24713)



The -intercept is the point on the graph where the -coordinate is zero. Substitute zero for and then do the arithmetic to determine the resulting value of . Then form the ordered pair where is the value of just calculated.

Use the same process to find the -intercept, except you substitute zero for .

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by lenny460(1073)

To find the x-intercept:
let y = 0
x - 2(0) = 0
x = 0

To find the y-intercept:
let x = 0
0 - 2y = 0
y = 0

The ordered pair is:

(0,0)


obuong3@aol.com